5^97 divided 52 the remainder obtained is a) 8 b) 1 c) 5 d) 4 e) 0 f) 3

Grade:12th pass

1 Answers

Arun
5 years ago

One method is to use modulo-arithmetic.

An example: To find (11^7) mod 13 do the following:

11^2 = 121 is congruent to 4(mod 13)

11^7 = ((11^2)^3) * 11

        => (4^3)*11 mod 13

        => (4^2)*4*11 mod 13

        => 16*4*11 mod 13

        =>  3*4*11 mod 13              Using: (a * b) mod c = ((a mod c )* (b mod c))mod c

        => 132 mod 13

        => 2 mod 13

i.e. 11^ 7 mod 13 is 2.

Using the same procedure, we get (5^97) mod 52 = 5

5^3 = 125 is congruent to 21 mod 52

(5^97) mod 52 => ((5^3)^32)*5 mod 52

                         =>  (21^32)*5  mod 52

                       Using 21^2 = 441 is congruent to 25 mod 52,

                        =>( (21^2)^16 )* 5 mod 52

                        => (25^16)* 5 mod 52

                        Using 25^2 = 625 is congruent to 1 mod 52,

                        => (1^8)* 5 mod 52

                        => 5 mod 52

i.e. the remainder when 5^97 is divided by 52 is 5.